# thanks: Corresponding Author (shared)thanks: These authors contributed equally to this workthanks: These authors contributed equally to this workthanks: Corresponding Author (shared)

### v1

@@ Abstract
The Kuramoto model and its generalizations have been broadly employed to characterize and mechanistically understand various collective dynamical phenomena, especially the emergence of synchrony among coupled oscillators. Despite almost five decades of research, many questions remain open, in particular for finite-size systems. Here, we generalize recent work [Phys. Rev. Lett. 130, 187201 (2023)] on the finite-size Kuramoto model with its state variables analytically continued to the complex domain and also complexify its system parameters. Intriguingly, systems of two units with purely imaginary coupling do not actively synchronize even for arbitrarily large magnitudes of the coupling strengths, $|K|\rightarrow\infty$, but exhibit conservative dynamics with asynchronous rotations or librations for all $|K|$. For generic complex coupling, both, traditional phase-locked states and asynchronous states generalize to complex locked states, fixed points off the real subspace that exist even for arbitrarily weak coupling. We analyze a new collective mode of rotations exhibiting finite, yet arbitrarily large winding numbers. Numerical simulations for large networks indicate a novel form of discontinuous phase transition. We close by pointing to a range of exciting questions for future research.
@@ I Introduction
Most results on the Kuramoto models have been obtained in the thermodynamic limit $N\rightarrow\infty$, where phase transitions occur in the form of bifurcations of lower-dimensional descriptions Ott and Antonsen (2008, 2009); Ott, Hunt, and Antonsen (2011); Marvel, Mirollo, and Strogatz (2009); Strogatz and Mirollo (1991). However, for finite-$N$ models despite them more closely describing reality, most results are purely numerical, sometimes employing intricate technical approaches Peter and Pikovsky (2018); Peter, Gong, and Pikovsky (2019) and analytical approaches that may offer stronger mechanistic insights are largely missing to date. Recent work Thümler et al. (2023) has studied the Kuramoto model analytically continued to complex state variables, in part to access finite-$N$ systems mathematically. The specific work uncovered a new form of synchrony and especially highlighted the option of understanding finite systems via analytically accessing complex locked states, fixed points that exist off the real invariant subspace in the complex domain even for small $K$. The complex locked states represent generalized forms of synchrony reached even if the coupling is not strong enough to induce synchrony in the real-variable system. In contrast to previous studies on finite systems, the existence of these fixed points enables analytic access to the dynamical states for finite systems without requiring standard phase-locking (and thus sufficiently strong coupling).
@@
In this article, we generalize the original real-variable Kuramoto model and analytically continue it to both complex state variablesThümler et al. (2023) and complex parameters. Our findings include conditions underlying attracting and non-attracting synchronous states, the coexistence of librating and rotating motions and motion with finite, yet arbitrarily large winding numbers, phenomena akin to the undamped and damped physical pendulum Strogatz (2000b). Numerical analysis also suggests a novel form of discontinuous phase transition at low coupling magnitudes.
@@ II.1 The original, real-variable system
Let us first consider a system of (real-valued) Kuramoto oscillators that characterizes the collective dynamics of a broad class of coupled phase oscillators Acebrón et al. (2005); Rodrigues et al. (2016). Such an oscillator is fully described by its phase variable $x_{n}$ that is either defined as a real number or as an element of the 1-torus, i.e. $x_{n}\in\mathbb{T}:=\mathbb{R}/2\pi\mathbb{Z}$ on the unit circle. Their dynamics is governed by
@@
for $n\in[N]:=\{1,2,\cdots,N\}$. In Eq. (1), the intrinsic local dynamics of each oscillator is determined by its natural (intrinsic) frequency $\omega_{n}\in\mathbb{R}$, often drawn from a given probability distribution. The broadness of the natural frequency distribution characterizes the degree of mismatch or heterogeneity of the oscillators. Also, the Kuramoto oscillators are diffusively coupled via a sinusoidal function with a real coupling strength $K\in\mathbb{R}$. For a positive $K>0$, the coupling between oscillators is attractive, tending to adjust their states and frequencies. For negative $K<0$, the coupling is repulsive such that oscillators tend to be repelled from each other. The collective dynamics of the system of coupled Kuramoto oscillators emerges from an interplay between the heterogeneity in the oscillators’ individual intrinsic frequencies $\omega_{n}$ and the coupling strength., cf. Sec. III.1.
@@ III.1 Two Real-valued Kuramoto Oscillators: Revisited
Let us first briefly revisit two coupled original (real-variable) Kuramoto oscillators. The system (1) has a phase-shift invariance since oscillators are diffusively coupled. Thus, in Eq. (1) with $N=2$, the phase difference $\Delta x:=x_{2}-x_{1}$ follows
@@
III.2 Two Complexfied Kuramoto Oscillators
Let us now study a system of two complexified oscillators, i.e., $N=2$ in Eq. (3). The anti-symmetric state-difference (diffusive) coupling reduces (3) to one ordinary differential equation for the difference $\Delta z:=z_{2}-z_{1}$ of the two complex variables. Taking the parameters $K=|K|e^{\textrm{i}\alpha}$ and defining $\Delta\omega:=\omega_{2}-\omega_{1}=|\Delta\omega|e^{\textrm{i}\gamma}\in\mathbb{C}$ with $\gamma:=\arg(\Delta\omega)\in\mathbb{R}$, the dynamics follows
@@ III.2 Two Complexfied Kuramoto Oscillators
The dynamics (9) is controlled by three (scalar, real) effective parameters: $c$, $\gamma$, and $\alpha$. In the following, we analyze the influence of each parameter on the system of two complexified oscillators, respectively.
@@ III.3 Real Natural Frequencies and Complex Coupling
First, consider a system of two complexified oscillators with real natural frequencies ($\gamma=0$ in Eq. (9)). The governing equation then reads
@@ III.3.1 Purely Imaginary Coupling: $\alpha=\frac{\pi}{2}$
Figure 2: Libration and rotation for purely imaginary coupling ($N=2$). (a) A phase portrait around three fixed points. Black solid dots indicate fixed points (16). The orange solid curve indicates a rotation (with the winding number $q=\infty$) whereas the red curves display librations ($q=0$) encircling antiphase fixed points and the purple curve indicates a libration around an in-phase synchrony fixed point. (b) The same fixed points as in (a) with separatrices (white solid curves) and the conserved quantity defined in Eq. (17). (c) A trajectory that rotates along the cylindrical axis with an infinite winding number $q=\infty$. All subfigures are obtained for $c=1$.
@@
Thus, both eigenvalues (15) are purely imaginary for all fixed points such that all of them are neutrally stable (centers). We emphasize that all the fixed points $\Delta z^{*}_{2k}$ with even indices represent complete in-phase synchrony with respect to their real parts, i.e., $\Delta x_{2k}=0$ (mod $2\pi$) whereas those with odd indices, $\Delta z^{*}_{2k+1}$, represent exact antiphase or splay states with $\Delta x_{2k+1}=\pi$ (mod $2\pi$). Thus, in contrast to the original, real-valued model, the complexified system possesses (neutrally stable) states exhibiting complete in-phase synchrony even though the intrinsic frequencies are different, $\Delta\omega\neq 0$ and the coupling strength is finite, $|K|<\infty$.
@@
Figure 3: Absence of transition to locking ($N=2$). Comparison between the Kuramoto model with panels a),b), purely imaginary coupling ($\alpha=\frac{\pi}{2}$) and c),d), the real-valued Kuramoto model ($\alpha=0$). (a) Time evolution of the magnitude $r(t)$ of the complex Kuramoto order parameter for a rotation (orange) and a libration around a locked state: in-phase synchrony $\Delta z^{*}_{0}$ (purple) and an antiphase state $\Delta z^{*}_{1}$ (red). Here, $c=1$. Inset: the complex Kuramoto order parameter (2) in the complex plane. (b) Time-averaged $r(t)\right>_{t}$ as a function of the coupling strength $|K|$. (c) Time evolution $r(t)$ for $c>1$ (no locked state) and $c<1$ (locked state). (d) Time-averaged $r(t)\right>_{t}$ as a function of the coupling strength $K>0$ for $c<1$. For all panels, we set $\Delta\omega=1$ and vary $c$ via $|K|$.
@@
is positive-definite. Thus, the fixed points of in-phase synchrony $(\Delta x_{2k}^{*},\Delta y_{2k}^{*})$ are local minima of $E(\Delta x,\Delta y)$ whereas fixed points of antiphase states $(\Delta x_{2k+1}^{*},\Delta y_{2k+1}^{*})$ are local maxima. As Fig. 2 (b) illustrates, separatrices in state space separate trajectories with librating motion and thus winding number $q=0$ from those with rotating motion and winding number $q=\infty$, see also Fig. 2 (c).
@@ III.3.2 Generic Complex Coupling: $\alpha\in(0,\frac{\pi}{2})$
Figure 4: Full range of integer winding numbers $q\in\mathbb{N}$ for generic complex coupling ($N=2$). (a) A trajectory (black solid curve) for $c=1$ and $\alpha=1.46004$ is depicted that circles around the cylinder with a finite winding number $q=5$ and eventually stops at a stable fixed point. Red dots indicate fixed points: stable (solid) in Eq. (22) and unstable (open) in Eq. (23). (b) The analytical form (solid curves) of the winding number $q$ is depicted as a function of the parameter $\alpha$, together with numerical results (solid dots). (c) The asymptotic behavior (dashed line, Eq. (25)) of the winding number $q$ as $\alpha\rightarrow\frac{\pi}{2}^{-}$: blue ($c=0.5<1$), red ($c=1$) and green ($c=1.5>1$).
The two-unit system (12) with $\alpha\in(0,\frac{\pi}{2})$ possesses stable fixed points
@@
for $k\in\mathbb{Z}$ that represent near in-phase states in their real parts, $\Delta x^{*}_{2k}\approx 0$ (mod $2\pi$) or $\Delta x^{*}_{2k}=c\cos\alpha+\mathcal{O}(c^{3})$ (mod $2\pi$) as $c\rightarrow 0$. In addition, the other fixed points, representing (near-) anti-phase states, are given as
@@
for $k\in\mathbb{Z}$. Linear stability analysis confirms that the locked states (22) are indeed linearly stable whereas the fixed points (23) are unstable. A bifurcation takes place at $\alpha=\frac{\pi}{2}$ wherein neutrally stable fixed points associated with in-phase synchrony are stabilized. Indeed, at a given fixed point $\Delta z^{*}$, the Jacobian has eigenvalues $f^{\prime}(\Delta z^{*})=-e^{\textrm{i}\alpha}\cos\big{(}\sin^{-1}(ce^{-\textrm{i}\alpha})\big{)}$ for $\Delta z^{*}=\sin^{-1}(ce^{-\textrm{i}\alpha})$. Their real parts become zero if $\text{Im}[e^{2\textrm{i}\alpha}-c^{2}]=\sin 2\alpha=0$, i.e. for $\alpha=0$ or $\frac{\pi}{2}$ and integer multiples thereof. For more details, see below Eq. (30).
@@
For a complex coupling with $\alpha\in(0,\frac{\pi}{2})$, the system possesses stable locked states (22). Moreover, a trajectory from a random initial condition does not exhibit a ceaseless rotation nor a libration around a fixed point, but rather it eventually reaches one of the stable equilibria after rotating around the cylindrical axis, i.e., with a finite winding number $q<\infty$. To see this, consider a trajectory starting from an initial condition that reads $\Delta x(0)=\Delta x_{2k}^{*}$ and $\Delta y(0)\gg 1$ for a given $k$. In Fig. 4 (a), a trajectory initiated from $\Delta x(0)=\Delta x_{-2}^{*}$ and $\Delta y(0)\gg 1$ is depicted (black solid curve) for $c=1$ and $\alpha<\frac{\pi}{2}$. The trajectory winds around the cylinder and eventually spirals down to a stable locked state, i.e., $(\Delta x_{4}^{*},\Delta y_{4}^{*})$. Thus, the specific trajectory takes the finite winding number $q=5$. We emphasize that the winding number of such a trajectory decreases down to (any) finite integer $q<\infty$ from $q=\infty$ as the phase-lag parameter decreases from $\alpha=\frac{\pi}{2}$ (purely imaginary coupling; $K\in\textrm{i}\mathbb{R}$) to $\alpha<\frac{\pi}{2}$ (complex coupling; $K\in\mathbb{C}\setminus\textrm{i}\mathbb{R}$).
@@
Recall that from a numerical integration for $\alpha=\frac{\pi}{2}$, e.g., in Fig. 2 (c), a trajectory around the $\Delta x$-axis exhibits a nearly sinusoidal curve with the amplitude $|\Delta y^{*}|=\sinh^{-1}c$. Also, a numerical integration for $\alpha\lesssim\frac{\pi}{2}$ with the initial condition described above shows a tilted sinusoidal trajectory near the $\Delta x$-axis (see Fig. 4 (a)). Thus, we assume the intermediate trajectory near the horizontal axis satisfies $\Delta y=-|\Delta y^{*}|\cos(\Delta x-\Delta x^{*})+A\Delta x=-\sinh^{-1}(c)\cos(\Delta x-\Delta x^{*})+A\Delta x$ with a slope $A$ and shifted horizontally by $\Delta x^{*}$. Also, assuming that the trajectory settles down to a stable fixed point $(\Delta x^{*}+2\pi k,\Delta y^{*})$ without spiraling down to it leads to $\Delta y^{*}=-\sinh^{-1}c+A\Delta x^{*}$. Then, the slope $A$ determines the winding number via $A=\frac{\Delta y^{*}}{2\pi q}=\frac{\Delta y^{*}+\sinh^{-1}c}{\Delta x^{*}}$. Solving this algebraic equation, an analytical form of the winding number as a function of parameters reads
@@
for $\alpha<\frac{\pi}{2}$. In Fig. 4 (b), the analytic form (24) of the winding number is depicted together with numerically obtained results for three different values of $c$. The actual winding number is an integer and thus a step function of $\alpha$. For the purpose of illustration, we therefore plot it (only) for the phase-lag parameter $\alpha$ that is numerically closest to the analytic curve. The good agreement between the numerical results and analytical prediction means that our analytical estimate (24) is very close to the actual winding number step function. We remark that our prediction (24) relies on the approximation of the rotating motion by an exact (tilted) sinusoidal function $\Delta y(\Delta x)$.
We underline that the winding number $q(\alpha)$ in Eq. (24) diverges as $\alpha\rightarrow\frac{\pi}{2}^{-}$, consistent with the findings in the system featuring a purely imaginary coupling, i.e., $\alpha=\frac{\pi}{2}$ where $q=\infty$ in Fig. 2 (c). Employing an asymptotic series expansion Bender and Orszag (1999) with respect to orders of $\frac{\pi}{2}-\alpha$, the leading order of the asymptotic behavior of the winding number reads
@@
as $\alpha\rightarrow\frac{\pi}{2}^{-}$. In Fig. 4 (c), we show the comparison of the analytic form of the winding number (24) with Eq. (25) near $\alpha=\frac{\pi}{2}$ from below. For any given $c$, the asymptotic behavior of the winding number follows Eq. (25) below near the purely imaginary coupling.

### v2

@@ Abstract
The Kuramoto model and its generalizations have been broadly employed to characterize and mechanistically understand various collective dynamical phenomena, especially the emergence of synchrony among coupled oscillators. Despite almost five decades of research, many questions remain open, in particular for finite-size systems. Here, we generalize recent work [Phys. Rev. Lett. 130, 187201 (2023)] on the finite-size Kuramoto model with its state variables analytically continued to the complex domain and also complexify its system parameters. Intriguingly, systems of two units with purely imaginary coupling do not actively synchronize even for arbitrarily large magnitudes of the coupling strengths, $|K|\rightarrow\infty$, but exhibit conservative dynamics with asynchronous rotations or librations for all $|K|$. For generic complex coupling, both, traditional phase-locked states and asynchronous states generalize to complex locked states, fixed points off the real subspace that exist even for arbitrarily weak coupling. We analyze a new collective mode of rotations exhibiting finite, yet arbitrarily large rotation numbers. Numerical simulations for large networks indicate a novel form of discontinuous phase transition. We close by pointing to a range of exciting questions for future research.
@@ I Introduction
Most results on the Kuramoto models have been obtained in the thermodynamic limit $N\rightarrow\infty$, where phase transitions occur in the form of bifurcations of lower-dimensional descriptions Ott and Antonsen (2008, 2009); Ott, Hunt, and Antonsen (2011); Marvel, Mirollo, and Strogatz (2009); Strogatz and Mirollo (1991). However, for finite-$N$ models despite them more closely describing reality, most results are purely numerical, sometimes employing intricate technical approaches Peter and Pikovsky (2018); Peter, Gong, and Pikovsky (2019) and analytical approaches that may offer stronger mechanistic insights are largely missing to date. Recent work Thümler et al. (2023) has studied the Kuramoto model analytically continued to complex state variables, yet still with all parameters staying real, in part to access finite-$N$ systems mathematically. The specific work uncovered a new form of synchrony and especially highlighted the option of understanding finite systems via analytically accessing complex locked states, fixed points that exist off the real invariant subspace in the complex domain even for small $K$. The complex locked states represent generalized forms of synchrony reached even if the coupling is not strong enough to induce synchrony in the real-variable system. In contrast to previous studies on finite systems, the existence of these fixed points enables analytic access to the dynamical states for finite systems without requiring standard phase-locking (and thus sufficiently strong coupling).
@@
In this article, we generalize the original real-variable Kuramoto model and analytically continue it to both complex state variablesThümler et al. (2023) and complex parameters. Our findings for $N=2$ coupled units include conditions underlying attracting and non-attracting synchronous states, the coexistence of librating and rotating motions and motion with finite, yet arbitrarily large rotation numbers, phenomena akin to the undamped and damped physical pendulum Strogatz (2000b). Numerical analysis for networks of $N=128$ also suggests a novel form of discontinuous phase transition at low coupling magnitudes.
@@ II.1 The original, real-variable system
Let us first consider a system of (real-valued) Kuramoto oscillators that characterizes the collective dynamics of a broad class of coupled phase oscillators Acebrón et al. (2005); Rodrigues et al. (2016). Such an oscillator is fully described by its phase variable $x_{n}$ that is either defined as a real number or as an angular coordinate of the 1-torus, i.e. $x_{n}\in\mathbb{T}:=\mathbb{R}/2\pi\mathbb{Z}$ on the unit circle. Their dynamics is governed by
@@
for $n\in[N]:=\{1,2,\cdots,N\}$. In Eq. (1), the intrinsic local dynamics of each oscillator is determined by its natural (intrinsic) frequency $\omega_{n}\in\mathbb{R}$, often drawn from a given probability distribution. The broadness of the natural frequency distribution characterizes the degree of mismatch or heterogeneity of the oscillators. Also, the Kuramoto oscillators are coupled via a sinusoidal function of phase difference with a real coupling strength $K\in\mathbb{R}$. For a positive $K>0$, the coupling between oscillators is attractive, tending to adjust their states and frequencies. For negative $K<0$, the coupling is repulsive such that oscillators tend to be repelled from each other. The collective dynamics of the system of coupled Kuramoto oscillators emerges from an interplay between the heterogeneity in the oscillators’ individual intrinsic frequencies $\omega_{n}$ and the coupling strength., cf. Sec. III.1.
@@ III.1 Two Real-valued Kuramoto Oscillators: Revisited
Let us first briefly revisit two coupled original (real-variable) Kuramoto oscillators. The system (1) has a phase-shift invariance since oscillators are coupled through their phase differences. Thus, in Eq. (1) with $N=2$, the phase difference $\Delta x:=x_{2}-x_{1}$ follows
@@
III.2 Two Complexified Kuramoto Oscillators
Let us now study a system of two complexified oscillators, i.e., $N=2$ in Eq. (3). The anti-symmetric state-difference coupling reduces (3) to one ordinary differential equation for the difference $\Delta z:=z_{2}-z_{1}$ of the two complex variables. Taking the parameters $K=|K|e^{\textrm{i}\alpha}$ and defining $\Delta\omega:=\omega_{2}-\omega_{1}=|\Delta\omega|e^{\textrm{i}\gamma}\in\mathbb{C}$ with $\gamma:=\arg(\Delta\omega)\in\mathbb{R}$, the dynamics follows
@@ III.2 Two Complexified Kuramoto Oscillators
The dynamics (9) is controlled by three (scalar, real) effective parameters: $c$, $\gamma$, and $\alpha$. In the following, we analyze the influence of each parameter on the system of two complexified oscillators.
@@ III.3 Real Natural Frequencies and Complex Coupling
First, consider a system of two complexified oscillators with real natural frequencies ($\omega_{1},\omega_{2}\in\mathbb{R}$ and so $\gamma=0$ in Eq. (9)). The governing equation then reads
@@ III.3.1 Purely Imaginary Coupling: $\alpha=\frac{\pi}{2}$
Figure 2: Libration and rotation for purely imaginary coupling ($N=2$). (a) A phase portrait around three fixed points. Black solid dots indicate fixed points (16). The orange solid curve indicates a rotation (with the rotation number $q=\infty$) whereas the red curves display librations ($q=0$) encircling antiphase fixed points and the purple curve indicates a libration around an in-phase synchrony fixed point. (b) The same fixed points as in (a) with separatrices (white solid curves) and the conserved quantity defined in Eq. (17). (c) A trajectory that rotates along the cylindrical axis with an infinite rotation number $q=\infty$. All subfigures are obtained for $c=1$.
@@
Thus, both eigenvalues (15) are purely imaginary for all fixed points so that all of them are neutrally stable (centers). We emphasize that all the fixed points $\Delta z^{*}_{2k}$ with even indices represent complete in-phase synchrony with respect to their real parts, i.e., $\Delta x_{2k}=0$ (mod $2\pi$) whereas those with odd indices, $\Delta z^{*}_{2k+1}$, represent exact antiphase or splay states with $\Delta x_{2k+1}=\pi$ (mod $2\pi$). Thus, in contrast to the original, real-valued model, the complexified system possesses (neutrally stable) states exhibiting complete in-phase synchrony even though the intrinsic frequencies are different, $\Delta\omega\neq 0$ and the coupling strength is finite, $|K|<\infty$.
@@
Figure 3: Absence of transition to locking ($N=2$). Comparison between the Kuramoto model with panels a),b), purely imaginary coupling ($\alpha=\frac{\pi}{2}$) and c),d), the real-valued Kuramoto model ($\alpha=0$). (a) Time evolution of the magnitude $r(t)$ of the complex Kuramoto order parameter for a rotation (orange) and a libration around a locked state: in-phase synchrony $\Delta z^{*}_{0}$ (purple) and an antiphase state $\Delta z^{*}_{1}$ (red). Here, $c=1$. Inset: the complex Kuramoto order parameter (2) in the complex plane. (b) Time-averaged $r(t)\right>_{t}$ as a function of the coupling strength $|K|$. For every $c$, each data point is obtained from an initial condition chosen close to the respective type of solution (libration around one or another fixed point or rotation). (c) Time evolution $r(t)$ for $c>1$ (no locked state) and $c<1$ (locked state). (d) Time-averaged $r(t)\right>_{t}$ as a function of the coupling strength $K>0$ for $c<1$. Data points are obtained from initial conditions randomly drawn from $\mathbb{T}^{2}$. For all panels, we set $\Delta\omega=1$ and vary $c$ via $|K|$.
@@
is positive-definite. Thus, the fixed points of in-phase synchrony $(\Delta x_{2k}^{*},\Delta y_{2k}^{*})$ are local minima of $E(\Delta x,\Delta y)$ whereas fixed points of antiphase states $(\Delta x_{2k+1}^{*},\Delta y_{2k+1}^{*})$ are local maxima. As Fig. 2 (b) illustrates, separatrices in state space separate trajectories with librating motion and thus rotation number $q=0$ from those with rotating motion and rotation number $q=\infty$, see also Fig. 2 (c). Here, the rotation number is defined as the integer number of turns that the rotational trajectory takes along the $\Delta x$-axis around the cylinder, $\Delta x\in\mathbb{T}$, i.e., the integer multiple
$q=\lim_{t\rightarrow\infty}\bigg{\lfloor}\frac{\Delta x(t)-\Delta x(0)}{2\pi}\bigg{\rfloor}$
of $2\pi$ that the variable $\Delta x$ increases along the trajectory. Here $\lfloor$.$\rfloor$ . ⌋ is the floor function.
@@ III.3.2 Generic Complex Coupling: $\alpha\in(0,\frac{\pi}{2})$
Figure 4: Full range of integer rotation numbers $q\in\mathbb{N}$ for generic complex coupling ($N=2$). (a) A trajectory (black solid curve) for $c=1$ and $\alpha=1.46004$ is depicted that circles around the cylinder with a finite rotation number $q=5$ and eventually stops at a stable fixed point. Red dots indicate fixed points: stable (solid) in Eq. (22) and unstable (open) in Eq. (23). (b) The analytical form (solid curves) of the rotation number $q$ is depicted as a function of the parameter $\alpha$, together with numerical results (solid dots). (c) The asymptotic behavior (dashed line, Eq. (25)) of the rotation number $q$ as $\alpha\rightarrow\frac{\pi}{2}^{-}$: blue ($c=0.5<1$), red ($c=1$) and green ($c=1.5>1$).
The two-unit system (12) with $\alpha\in(0,\frac{\pi}{2})$ possesses fixed points
@@
for $k\in\mathbb{Z}$ that represent near in-phase states in their real parts ($\Delta x^{*}_{2k}=0$), in particular $\Delta x^{*}_{2k}=c\cos\alpha+\mathcal{O}(c^{3})$ (mod $2\pi$) asymptotically as $c\rightarrow 0$. In addition, the other fixed points, representing (near-) anti-phase states, are given as
@@
for $k\in\mathbb{Z}$. Linear stability analysis from (15) shows that the complex locked states (22) are linearly stable whereas the states (23) are unstable. A bifurcation takes place at $\alpha=\frac{\pi}{2}$ wherein neutrally stable fixed points associated with in-phase synchrony are stabilized. Indeed, at a given fixed point $\Delta z^{*}$, the Jacobian has eigenvalues $f^{\prime}(\Delta z^{*})=-e^{\textrm{i}\alpha}\cos\big{(}\sin^{-1}(ce^{-\textrm{i}\alpha})\big{)}$ for $\Delta z^{*}=\sin^{-1}(ce^{-\textrm{i}\alpha})$. Their real parts become zero if $\text{Im}[e^{2\textrm{i}\alpha}-c^{2}]=\sin 2\alpha=0$, i.e. for $\alpha=0$ or $\frac{\pi}{2}$ and integer multiples thereof. For more details, see below Eq. (30).
@@
For a complex coupling with $\alpha\in(0,\frac{\pi}{2})$, the system possesses stable locked states (22). Moreover, a trajectory from a random initial condition does not exhibit a ceaseless rotation nor a libration around a fixed point, but rather it eventually reaches one of the stable equilibria after rotating around the cylindrical axis, i.e., with a finite rotation number $q<\infty$. To see this, consider a trajectory starting from an initial condition that reads $\Delta x(0)=\Delta x_{2k}^{*}$ and $\Delta y(0)\gg 1$ for a given $k$. In Fig. 4 (a), a trajectory initiated from $\Delta x(0)=\Delta x_{-2}^{*}$ and $\Delta y(0)\gg 1$ is depicted (black solid curve) for $c=1$ and $\alpha<\frac{\pi}{2}$. The trajectory winds around the cylinder and eventually spirals down to a stable locked state, i.e., $(\Delta x_{4}^{*},\Delta y_{4}^{*})$. Thus, the specific trajectory takes the finite rotation number $q=5$. We emphasize that the rotation number of such a trajectory decreases down to (any) finite integer $q<\infty$ from $q=\infty$ as the phase-lag parameter decreases from $\alpha=\frac{\pi}{2}$ (purely imaginary coupling; $K\in\textrm{i}\mathbb{R}$) to $\alpha<\frac{\pi}{2}$ (complex coupling; $K\in\mathbb{C}\setminus\textrm{i}\mathbb{R}$).
@@
Recall that from a numerical integration for $\alpha=\frac{\pi}{2}$, e.g., in Fig. 2 (c), a trajectory around the $\Delta x$-axis exhibits a nearly sinusoidal curve with the amplitude $|\Delta y^{*}|=\sinh^{-1}c$. Also, a numerical integration for $\alpha\lesssim\frac{\pi}{2}$ with the initial condition described above shows a tilted sinusoidal trajectory near the $\Delta x$-axis (see Fig. 4 (a)). Thus, we assume the intermediate trajectory near the horizontal axis satisfies $\Delta y=-|\Delta y^{*}|\cos(\Delta x-\Delta x^{*})+A\Delta x=-\sinh^{-1}(c)\cos(\Delta x-\Delta x^{*})+A\Delta x$ with a slope $A$ and shifted horizontally by $\Delta x^{*}$. Also, assuming that the trajectory settles down to a stable fixed point $(\Delta x^{*}+2\pi k,\Delta y^{*})$ without spiraling down to it leads to $\Delta y^{*}=-\sinh^{-1}c+A\Delta x^{*}$. Then, the slope $A$ determines the rotation number via $A=\frac{\Delta y^{*}}{2\pi q}=\frac{\Delta y^{*}+\sinh^{-1}c}{\Delta x^{*}}$. Solving this algebraic equation, an analytical form of the rotation number as a function of parameters reads
@@
for $\alpha<\frac{\pi}{2}$. In Fig. 4 (b), the analytic form (24) of the rotation number is depicted together with numerically obtained results for three different values of $c$. The actual rotation number is an integer and thus a step function of $\alpha$. For the purpose of illustration, we therefore plot it (only) for the phase-lag parameter $\alpha$ that is numerically closest to the analytic curve. The good agreement between the numerical results and analytical prediction means that our analytical estimate (24) is very close to the actual rotation number step function. We remark that our prediction (24) relies on the approximation of the rotating motion by an exact (tilted) sinusoidal function $\Delta y(\Delta x)$.
We underline that the rotation number $q(\alpha)$ in Eq. (24) diverges as $\alpha\rightarrow\frac{\pi}{2}^{-}$, consistent with the findings in the system featuring a purely imaginary coupling, i.e., $\alpha=\frac{\pi}{2}$ where $q=\infty$ in Fig. 2 (c). Employing an asymptotic series expansion Bender and Orszag (1999) with respect to orders of $\frac{\pi}{2}-\alpha$, the leading order of the asymptotic behavior of the rotation number reads
@@
as $\alpha\rightarrow\frac{\pi}{2}^{-}$. In Fig. 4 (c), we show the comparison of the analytic form of the rotation number (24) with Eq. (25) near $\alpha=\frac{\pi}{2}$ from below. For any given $c$, the asymptotic behavior of the rotation number follows Eq. (25) for $\alpha\to\frac{\pi}{2}^{-}$, i.e., as the complex coupling approaches purely imaginary coupling.